# Slice-Level Scheduling for High Throughput and Load Balanced LLM Serving

### v1

@@ 1 Sort $\mathcal{R}$ by the request input length in ascending order
// States denoting the total serving time
@@
        // Separate a request into a batch
@@ 14            $t\leftarrow\mathcal{T}[j-1]$ + $T_{serve}(i-j+1,L_{i},S)$
                    // Update total serving time
@@
                  $\mathcal{P}[i]\leftarrow j-1$ // Record splitting position
@@ 25      $i\leftarrow p$
Pseudocode of the bacthing algorithm is presented in Algorithm 1. The batching algorithm first sorts requests in ascending order by their input lengths. Then the state array $\mathcal{T}$ is created, where $\mathcal{T}[i]$ denotes the minimal total batch serving time for the first $i$ requests. The initial state $\mathcal{T}[0]=0$ indicates that the total serving time is 0 when there is no request to batch. The position array $\mathcal{P}$ stores positions for partitioning requests into batches.
During dynamic programming, the algorithm sequentially traverses each request. For each request, the algorithm first treats it as a separate batch, and then tries to batch it together with preceding requests to minimize the total batch serving time while guaranteeing that OOM errors will not occur. The state transfer function of dynamic programming is expressed by
(10) $\mathcal{T}[i]=\min\limits_{0<j\leq i}(\mathcal{T}[j-1]+T_{serve}(i-j+1,L_{i},S)),$
where $i-j+1$ computes the batch size. Since requests are sorted by the request input length at the beginning, when traversing to the $i$th request, its input length is the maximal input length of the first $i$ requests, and hence its input length can be utilized as the batch input length $L_{i}$ to calculate the estimated batch serving time $T_{serve}$. After the dynamic programming, the requests are divided into batches according to the recorded splitting positions.
In the batching algorithm, whether OOM errors will occur is determined by the memory usage estimator, and the estimated serving time of batches is calculated by the serving time estimator.
4.5. Balanced Load-Oriented Offloading
The offloader achieves load balancing among multiple workers using a max-min-based algorithm. The load of a worker is defined as the time it takes to serve all the batches in its local queue and the initial load of each worker is 0.
@@ 4.5. Balanced Load-Oriented Offloading
(11) $T_{load}(w)\leftarrow T_{load}(w)+T_{serve}(N,L_{i},S),$
@@
In addition, after each time a worker completes batch serving, SCLS updates the load of the worker by subtracting the estimated serving time of the batch it just served from its load. This can prevent the estimation error of serving time from accumulating in the load, thus enhancing the load-balancing.
4.6. Adaptive Schedule Interval Update
SCLS periodically fetches requests from the request pool and hands them over to the adaptive batcher at a time interval of $T$. After each batch offloading, SCLS updates the time interval $T$ by
(12) $T\leftarrow\max(\lambda\cdot\min\limits_{w}(T_{load}(w)),\Gamma),$
@@ 4.6. Adaptive Schedule Interval Update
Accordingly, when the load of each worker is light, $T$ decreases, so that requests can be processed by workers in time without waiting for a long time in the request pool, thus reducing the request response time. When there are a lot of batches waiting in the local queue of each worker, $T$ increases so that the adaptive batcher is highly likely to batch more requests together, thus increasing the throughput. Due to the serving time estimation error, $\lambda$ is introduced to prevent workers from being idle when the load is over-estimated, and $\Gamma$ is introduced to avoid feeding too few requests to the adaptive batcher when the load is under-estimated.
@@
$L$leftarrow $L_{i}$ + $S$ // $L\leq 2048$ under experimental settings
@@
      return $N>28$ // Total token number $L\leq 512$
@@
5. Experimental Evaluation
5.1. Experiment Setup
Testbed. Experiments are conducted on our testbed equipped with 8 NVIDIA A100 80GB GPUs connected over NVLink, 800GB CPU memory, and two 2.90GHz 32-core Intel(R) Xeon(R) CPUs.
Settings. We set both the maximal raw request input length and the maximal generation length limit to 1024. When the raw input length of a request is greater than 1024, it is truncated to 1024. When the total number of generated tokens of a request is greater than 1024, the request is returned.
Implementation. We implement SCLS and workers for huggingface transformers (HF) v4.35.0 and deepspeed inference (DS) v0.13.3 using Python 3.8. For the serving time estimator, we use curve_fit function from the scipy library (virtanen2020fundamental, ) to fit Eq. (3) and Eq. (4) with profiled per-iteration latency data. The memory estimator leverages different methods for the two inference engines to judge whether OOM will occur. For HF, Eq. (9) is used where $\zeta$ is set to 0.9. For DS, under the experimental settings, it determines if OOM errors will occur by rules that are presented in Algorithm 2. To fetch more requests while keeping workers busy, SCLS dynamically adjusts the schedule time interval $T$ by Eq. (12), where $\lambda$ is set to 0.5, $\Gamma$ is set to 6s and 3s for HF and DS according to their inference speed, respectively. Moreover, the slice length is set to 128 if not specified.
Workflow. We deploy 8 LLaMA2-13B workers on 8 NVIDIA A100 80GB GPUs. After LLM instances are loaded, we send requests from the CodeFuse request trace in the order they actually arrived. Requests are sent for 10 minutes and the request arrival times are generated using Poisson distribution with various request rates.
Metrics. We use request throughput, average response time, and 95% tail response time as evaluation metrics, where the response time is the total time between the user sending the request and receiving the generated results.
@@ 5.1. Experiment Setup
Figure 12. Request throughput, average response time and tail response time under various arrival rates.
@@
Figure 13. Dive into the superiority of SCLS.
Baselines. We compare SCLS with SLS and ILS schedulers. Since HF only supports static batching, when using HF as the inference engine, we only compare SCLS with SLS. For DS, we not only compare SCLS with SLS but also use Deepspeed-Fastgen, an ILS scheduler implemented on top of DS, to compare with SCLS.
  • Sequence-level Scheduling (SLS) (tensorflowserving, ; triton, ): The scheduler offloads requests to workers using the round-robin policy. Workers serve received requests in an FCFS manner with a fixed batch size. Since different engines have various $M_{engine}$, for HF and DS workers, the fixed batch size is set to 16 and 12 to avoid OOM errors, respectively.
  • Iteration-level Scheduling (ILS) (deepspeed-fastgen, ): Deepspeed-FastGen, the ILS scheduler built on top of DS, is one of the state-of-the-art ILS schedulers. It achieves a low latency with continuous batching and split-fuse technologies. However, it utilizes the round-robin policy to offload requests and adopts a conservative memory management mechanism which limits the throughput. For DS, we take FastGen-v0.2.0 as the ILS scheduler to compare with SCLS.
@@
Figure 14. Dive into the overhead of SCLS.
5.2. Overall Performance
In this subsection, we compare the performance of SCLS to the baseline scheduling techniques under different request arrival rates.
@@ 5.2. Overall Performance
The experimental results are depicted in Fig. 12, where different colors represent different inference engines, and various hatch styles represent various scheduling techniques. We can see that SCLS always outperforms SLS and ILS and shows the best performance when integrating with various inference engines under different request arrival rates.
Compared with SLS, the superiority of SCLS mainly comes from three sources. The first is to return completed requests in time, thus avoiding request waiting and greatly reducing the number of generated invalid tokens. The second is to serve requests with a large batch size to achieve a high throughput. The third is to reduce the serving latency by mitigating the padding overhead using the adaptive batcher. Due to the different memory management mechanisms of HF and DS, the maximal batch size that can be supported under the same batch input length and the same slice length is different, and SCLS can obtain different performance gains when integrated with various inference engines.
HF can flexibly support variable batch sizes. Given the request input length and the slice length, SCLS’s adaptive batcher can batch more requests together to achieve a larger batch size when using HF as the inference engine. As shown in Fig. 12, SCLS can increase the throughput by 232.3% to 315.8%, decrease the average response time by 82.4% to 84.9%, and decrease the tail response time by 79.8% to 81.0% compared with SLS when integrated with HF.
DS’s memory management mechanism is not flexible enough to support variable batch sizes, and hence we use the conservative rules presented in Algorithm 2 to prevent OOM errors. Therefore, when SCLS is integrated with DS, the throughput is only increased by 82.5% to 191.9% compared with SLS, which is less than the gain when integrated with HF. However, due to the timely return of requests and low padding overhead, SCLS can still reduce the average response time by 80.8% to 91.1% and the tail response time by 76.6% to 88.9% compared with SLS when integrated with DS.
@@
Compared with ILS, SCLS can support larger batch sizes without causing OOM errors, whereas Deepspeed-FastGen adopts a conservative memory management policy to prevent OOM and speedup inference by limiting the number of parallel-processing requests. As a result, SCLS can improve the throughput by 61.6% to 171.0%, reduce the average request response time by 78.4% to 88.3%, and reduce the tail response time by 74.6% to 85.8% compared with ILS when integrated with DS.
@@
Figure 15. Request throughput, average response time and tail response time under various strategies.
@@
Figure 16. Dive into the results of ablation experiments.
To dive into the superior performance of SCLS, we present the average invalid token number of requests, the average batch size, and the average pad token number of requests for SLS and SCLS at various request rates when integrated with various inference engines in Fig. 13. For SCLS, if a request is rescheduled multiple times, we add up the number of pad tokens at each schedule as its pad token number. From Fig. 13(a), we can see that under SCLS, the number of invalid tokens generated in the request waiting process is greatly reduced, which implies that the completed request can be returned in time through generation slicing. Besides, we can also find that the batch size is greatly improved from Fig. 13(b). The improvement is more significant when SCLS is integrated with HF, which is 100.3% to 226.1%. For DS, the batch size improvement of SCLS over SLS is only 43.2% to 86.1%, which is not as much as integration with HF due to the inflexible memory management mechanism of DS. Moreover, as the request rate increases, more requests of varying input and generation lengths are sent to the adaptive batcher, which provides more opportunities for the batching algorithm to batch more requests together and reduce the computational overhead of pad tokens. Thus, from Fig. 13(b) and Fig. 13(c), we can see that under SCLS, the batch size increases with the request rate while the number of pad tokens decreases with the request rate.
Figure 17. Load imbalance under various request rates
The overhead of SCLS comes from the re-computation of the prefill phase with each reschedule. We illustrate the distribution of request reschedule number (a.k.a. slice number) under different request rates in Fig. 14(a). We find that the vast majority of requests have less than three reschedules, this is because their request generation lengths are small and can complete generation within three slices. Although the total serving time of requests with long generation lengths is increased due to multiple re-computations of the prefill phase, they also benefit from SCLS as their queuing time is greatly reduced due to increased throughput.
In Section 4.2, we mentioned that if all requests generate EOS before the generation length reaches the slice length, the requests will be returned early, which leads to inaccurate time estimation. To verify that this happens very rarely, in Fig. 14(b), we calculate the proportion of these early returned batches to the total number of served batches. We can find that early return happens rarely, with a maximal proportion less than 1%, and hence the early return has negligible effect on the performance of SCLS.
5.3. Load Balance
We calculated the standard deviation (STD) of each instance’s completion time (CT) at the end of the experiment in Fig. 17.
We find that the standard deviation of SCLS is always the smallest, which indicates that the completion time of each instance is similar and SCLS achieves load balancing across multiple LLM instances. This is because the offloading algorithm can accurately estimate the load of workers and leverage the max-min algorithm to reduce the load gap between workers.
@@ 5.3. Load Balance
Figure 18. Request throughput, average response time and tail response time under various slice lengths.
@@
Figure 19. Dive into the slice length impact on SCLS performance .
5.4. Ablation Studies
In this section, we validate the performance gain of the design features of SCLS. We keep adding features to SLS to constitute multiple strategies and then verify their performance in terms of request throughput and response time at a request arrival rate of 20. The experimental results are presented in Fig. 15.
@@ 5.4. Ablation Studies
Figure 20. Dive into the slice length impact on SCLS overhead.
Figure 21. Load imbalance under various slice lengths
5.5. Impact of Slice Length
From Fig. 18, we can find that with the increase in slice length, the performance first increases and then decreases. This is because when the slice length is set too small, there are a lot of reschedules of requests. The reschedules cause requests to be padded multiple times and hence the computational overhead caused by padding is severe when the slice length is set small as shown in Fig. 19(c). Besides, the frequent reschedules cause a large number of prefill re-computations. Although with a small slice length, completed requests can be returned in a timely manner and the batch size can be relatively large, it is not enough to make up for the computational overhead of the large amount of re-padding and prefill re-computations. When the slice length is increased, although the batch size decreases slightly, the number of reschedules decreases dramatically as shown in Fig. 20(a), and hence the performance is improved. However, setting the slice length too long not only further decreases the batch size as shown in Fig. 19(b), but also causes more serious request waiting, and more invalid tokens are generated as shown in Fig.19(a), so the performance decreases. Furthermore, a long slice length causes a large early return ratio as shown in Fig. 20(b), which makes the serving time estimation inaccurate, and hence the workload becomes unbalanced as shown in Fig. 21. Therefore, when setting the slice length, we need to do a trade-off between the request padding, the batch size, the prefill re-computations, and the request waiting.
@@ 5.5. Impact of Slice Length
Figure 22. Scalability of SCLS.
5.6. Scalability
@@
6. Related Works
Model compression. Quantization (frantar2023optq, ; xiao2023smoothquant, ; dettmers2022gpt3, ; shao2023omniquant, ; chee2024quip, ) reduces the numerical precision of model weights to lower-bit representations to reduce the memory consumed by model parameters. Knowledge distillation (liang2020mixkd, ; li2021dynamic, ; liu2022multi, ; zhu2023minigpt, ; xu2024wizardlm, ) trains small-scale student models using outputs of large teacher models to compress the model size while maintaining prediction accuracy. Pruning (xu2022dense, ; sun2023simple, ; liu2023deja, ) removes redundant network modules from LLMs and improves the inference speed. These works are orthogonal to SCLS, and the computational efficiency of SCLS can be further enhanced with compressed LLMs.
Speculative decoding. Speculative decoding (leviathan2023fast, ; xia2023speculative, ; sun2024spectr, ; kim2024speculative, ) leverages a small language model to autoregressively generate draft outputs and refine the draft with the large model periodically. However, the increase in serving efficiency of speculative decoding depends on the predictive power of the small model. Besides, most speculative decoding algorithms do not support batch processing and are not applicable to high-concurrency scenarios.

### v2

@@ 1 Sort $\mathcal{R}$ by the request input length in ascending order
// States denoting total serving time
@@
       /* Separate the $i$th request into a single batch */
@@ 14            $t\leftarrow\mathcal{T}[j-1]$ + $T_{serve}(i-j+1,L_{i},S)$
                   /* Update the total serving time */
@@
                  $\mathcal{P}[i]\leftarrow j-1$ /* Record the splitting position */
@@ 25      $i\leftarrow p$
IV-E Balanced Load-Oriented Offloading
The offloader achieves load balancing among multiple workers using a max-min-based algorithm. The load of a worker is defined as the time it takes to serve all the batches in its local queue, and the initial load of each worker is 0.
@@ IV-E Balanced Load-Oriented Offloading
$T_{load}(w)\leftarrow T_{load}(w)+T_{serve}(N,L_{i},S),$ (11)
@@
After each time a worker completes batch serving, SCLS subtracts the estimated serving time of the batch it just served from its load. This can prevent the estimation error of serving time from accumulating in the worker’s load.
IV-F Adaptive Schedule Interval Update
SCLS periodically fetches requests from the request pool and hands them over to the adaptive batcher at a time interval of $T$. After each batch offloading, SCLS updates $T$ by
$T\leftarrow\max(\lambda\cdot\min\limits_{w}(T_{load}(w)),\Gamma),$ (12)
@@ IV-F Adaptive Schedule Interval Update
When the load of each worker is light, $T$ decreases so that requests can be processed by workers in time without waiting for a long time in the request pool, thus reducing the request response time. When there are a lot of batches waiting in the local queue of each worker, $T$ increases so that the adaptive batcher is highly likely to batch more requests together, thus increasing the throughput. Due to the serving time estimation error, $\lambda$ is introduced to prevent workers from being idle when the load is over-estimated, and $\Gamma$ is introduced to avoid feeding too few requests to the adaptive batcher when the load is under-estimated.
@@
$L$leftarrow $L_{i}$ + $S$ /* $L\leq 2048$ under the experimental settings */
@@
      return $N>28$ /* Total token number $L\leq 512$/
@@
V Experimental Evaluation
V-A Experiment Setup
Experiments are conducted in the production environment at Ant Group, where 8 LLaMA-13B LLM instances are deployed with 8 NVIDIA A100 80GB GPUs connected over NVLink. We use LLaMA2-13B in the experiments because LLaMA’s model architecture is representative and has been adopted by many LLMs. Besides, 13B-parameter LLMs can meet the needs of most real-world application scenarios in the production environment. We take the request throughput, average and tail response time as metrics to evaluate the performance of SCLS, following existing work on LLM serving [4, 5, 14].
We implement SCLS, DS-worker and HF-worker with Python 3.8. For the serving time estimator, we use curve_fit function from the scipy library [15] to fit Eq. (3) and Eq. (4). For the memory usage estimator, $\zeta$ in Eq. (9) is set to 0.9 for HF. For DS, SCLS determines if OOM errors will occur by rules that are presented in Algorithm 2. To fetch more requests while keeping workers busy, SCLS dynamically adjusts the schedule time interval $T$ by Eq. (12), where $\lambda$ is set to 0.5, $\Gamma$ is set to 6s and 3s for HF and DS according to their inference speed, respectively. Moreover, the slice length is set to 128 if not specified.
Requests from CodeFuse’s traces are sent in the order they actually arrived for 10 minutes. Arrival times are generated using Poisson distribution with various request rates. We set both the maximal raw request input length and the maximal generation length limit to 1024, covering length range settings of existing studies [16, 17, 4, 5, 18].
We compare SCLS with SLS and ILS schedulers. Since HF only supports static batching, for HF, we only compare SCLS with SLS. For DS, we compare SCLS with SLS and Deepspeed-Fastgen, an ILS scheduler implemented on top of DS. We don’t compare SCLS with vLLM [19] because vLLM does not support static batching, and hence SCLS can not achieve the same kernel-level optimization with vLLM, which makes it challenging to achieve a fair comparison from a pure scheduling perspective. Besides, Deepspeed-Fastgen has been shown to outperform vLLM’s performance [6]. Therefore, we think it reasonable and convincing to compare SCLS with Deepspeed-Fastgen.
  • Sequence-level Scheduling (SLS) [3, 2]: The scheduler offloads requests to workers using the round-robin policy. Workers serve received requests in an FCFS manner with a fixed batch size. Since different engines have various $M_{engine}$, for HF and DS workers, the fixed batch size is set to 16 and 12 to avoid OOM errors, respectively.
  • Iteration-level Scheduling (ILS) [6]: Deepspeed-FastGen, the ILS scheduler built on top of DS, achieves a low latency with continuous batching [4], paged attention[5], and chunked-prefill [20] techniques. However, it utilizes the round-robin policy to offload requests and adopts a conservative memory management mechanism that limits the number of parallel-processing requests. Since Deepspeed-FastGen is also implemented based on DS, when SCLS is integrated with DS, SCLS and ILS have the same kernel-level optimization, so a fair comparison can be made from a pure scheduling perspective.
@@ V-A Experiment Setup
Figure 9: Request throughput, average response time and tail response time under various arrival rates.
@@
Figure 10: Dive into the superiority of SCLS.
@@
Figure 11: Dive into the overhead of SCLS.
V-B Overall Performance
@@
We compare the performance of SCLS to the baseline schedulers in terms of request throughput and tail response time. The experimental results are depicted in Fig. 9, where different colors represent different inference engines, and various hatch styles represent various scheduling techniques. We can see that SCLS always outperforms SLS and ILS when integrating with various inference engines under different request arrival rates, as shown in Fig. 9.
For HF, SCLS can increase the throughput by 232.3% to 315.8%, decrease the average response time by 82.4% to 84.9%, and decrease the tail response time by 79.8% to 81.0% compared with SLS. For DS, SCLS increases the throughput by 82.5% to 191.9%, reduces the average response time by 80.8% to 91.1%, and decreases the tail response time by 76.6% to 88.9% compared with SLS.
@@
Compared with ILS, SCLS can support larger batch sizes without causing OOM errors, whereas Deepspeed-FastGen adopts a conservative memory management policy by limiting the number of parallel-processing requests. As a result, SCLS can improve the throughput by 61.6% to 171.0%, reduce the average request response time by 78.4% to 88.3%, and reduce the tail response time by 74.6% to 85.8% compared with ILS when integrated with DS.
Compared with SLS, the superiority of SCLS mainly comes from three sources. The first is to return completed requests in time, thus avoiding request waiting and greatly reducing the number of generated invalid tokens, as shown in Fig. 10(a). The second is to serve requests with a large batch size to achieve high throughput, as shown in Fig. 10(b). The third is to reduce the serving latency by mitigating the padding overhead using the adaptive batcher as shown in 10(c). Due to the different memory management mechanisms of HF and DS, the maximal batch size that can be supported under the same batch input length and the same slice length is different, and therefore, SCLS obtains different performance gains when integrated with various inference engines.
V-C SCLS Overhead
The overhead of SCLS comes from the re-computation of the prefill phase with each reschedule. We illustrate the distribution of request reschedule number (a.k.a. slice number) under different request rates in Fig. 11(a). We find that the vast majority of requests have less than three reschedules. This is because their request generation lengths are small and can complete generation within three slices. Although the total serving time of requests with long generation lengths is increased due to multiple re-computations of the prefill phase, they also benefit from SCLS as their queuing time is greatly reduced due to increased throughput.
In Section IV-B, we mentioned that if all requests generate EOS before the generation length reaches the slice length, the requests will be returned early, which leads to inaccurate time estimation. To verify that this happens very rarely, in Fig. 11(b), we calculate the proportion of these early returned batches to the total number of served batches. We can find that early return happens rarely, with a maximal proportion less than 1%. Hence, the early return has a negligible effect on the performance of SCLS.
V-D Load Balance
We calculated the standard deviation (STD) of each instance’s completion time (CT) at the end of the experiment in Fig. 12. We find that the CT STD of SCLS is always the smallest, which indicates that the completion time of each instance is similar and SCLS achieves load balancing across multiple LLM instances. This is because the offloading algorithm can accurately estimate the load of workers and leverage the max-min algorithm to reduce the load gap between workers.
Figure 12: Load imbalance under various request rates
@@ V-D Load Balance
Figure 13: Request throughput, average response time and tail response time under various strategies.
@@
Figure 14: Dive into the results of ablation experiments.
V-E Ablation Studies
We demonstrate the performance gain of the design features of SCLS. We keep adding features to SLS to constitute multiple strategies and then verify their performance in terms of request throughput and response time at a request arrival rate of 20. The experimental results are presented in Fig. 13.
@@ V-E Ablation Studies
Figure 15: Request throughput, average response time and tail response time under various slice lengths.
V-F Impact of Slice Length
From Fig. 15, we can find that with the increase in slice length, the performance first increases and then decreases. This is because when the slice length is set too small, there are a lot of reschedules of requests. The reschedules cause requests to be padded multiple times, and hence the computational overhead caused by padding is severe when the slice length is set small, as shown in Fig. 16(c). Besides, the frequent reschedules cause a large number of prefill re-computations. Although with a small slice length, completed requests can be returned in a timely manner, and the batch size can be relatively large, it is not enough to make up for the computational overhead of the large amount of re-padding and prefill re-computations. When the slice length is increased, although the batch size decreases slightly, the number of reschedules decreases dramatically as shown in Fig. 17(a), and hence the performance is improved. However, setting the slice length too long not only further decreases the batch size as shown in Fig. 16(b), but also causes more serious request waiting, and more invalid tokens are generated as shown in Fig.16(a), so the performance decreases. Furthermore, a long slice length causes a large early return ratio as shown in Fig. 17(b), which makes the serving time estimation inaccurate, and hence the workload becomes unbalanced as shown in Fig. 18. Therefore, when setting the slice length, we need to do a trade-off between the request padding, the batch size, the prefill re-computations, and the request waiting.
Since the performance of SCLS monotonically increases first and then monotonically decreases with the increase of the slice length, for application scenarios with specific request input and output distribution, we can simply use the binary search to quickly find the optimal setting of slice length with a time complexity of $\log(n)$, where $n$ is the possible maximal slice length, making the search for the optimal slice length simple and efficient.
@@ V-F Impact of Slice Length
Figure 16: Dive into the slice length impact on SCLS performance .
@@
Figure 17: Dive into the slice length impact on SCLS overhead.
Figure 18: Load imbalance under various slice lengths
V-G Handling Long-Generation Requests
We increase the proportion of requests with generation length greater than 512 to 60% and re-conduct the experiments at a request rate of 8. Since ILS has been demonstrated to be consistently superior to SLS, we only compare SCLS with ILS using DS as the inference engine. The results show that SCLS can improve the request throughput by 117% compared with the ILS scheduler. Besides, the average response time is reduced by 15.7% when the slice length increases from 128 to 256 due to greatly decreased re-prefill, which confirms that SCLS can effectively handle long generation scenarios by properly increasing the slice length.
@@ V-G Handling Long-Generation Requests
Figure 19: Scalability of SCLS.
V-H Scalability
@@
VI Related Works
VI-A Model Compression
. Model quantization [21, 22, 23, 24, 25] converts the numerical weight precision of LLMs into fewer bits to reduce the memory usage of LLM parameters. The technique of pruning [26, 27, 28, 29] eliminates redundant components of LLMs to increase the inference speed. These approaches can affect LLM accuracy while SCLS is a request scheduling framework that does not make any modification to LLM parameters and does not affect LLM accuracy. Besides, since SCLS is an inter-instance request scheduling framework, these works are orthogonal to SCLS, and SCLS can also improve the serving efficiency for quantized and pruned LLMs.
VI-B Low-level Optimization.
Kernel optimization techniques such as flash attention [30, 31] and paged attention [5] speed up LLM inference by implementing efficient CUDA kernels to improve computational efficiency. Parallelism optimization [32, 33, 34, 35, 36] proposes to split the model at intra-layer and inter-layer granularity to enable distributed LLM inference. These studies are always employed to make a single LLM instance more efficient in inference. However, SCLS is a scheduling system that schedules requests to multiple workers. It leverages proposed adaptive batching and scheduling algorithms to fully use the inference engine and balance load among multiple workers. Moreover, these studies are orthogonal to SCLS as well. SCLS can be augmented with these low-level techniques.
VI-C LLM Serving Systems.